# Bridging Training and Execution via Dynamic Directed Graph-Based Communication in Cooperative Multi-Agent Systems

### v1

@@ 4.2 Graph Coarsening Network
We constructed the graph coarsening network based on Equation (2) to address the last of the five communication problems outlined in Section 1. The schematics of the graph coarsening network are shown in Figure 2(a). The dynamic directed graph, composed of multiple subgraphs, can be considered a coarsened graph, where each subgraph is a super node to fit the global state. This approach enables hierarchical learning of global information. According to the theoretical analysis in Section 4.1, the input of the graph coarsening network consists of each agent’s local observations $o^{i}_{t}$, the adjacency trajectory matrix $A^{i}_{t}$ and the received message $\tilde{m}^{i}_{t}$. We normalize the adjacency trajectory matrix to prevent degree bias and gradient vanishing during training. We define $[\tilde{m}^{i}_{t}\parallel o^{i}_{t}]$ as $X^{i}_{t}$ as feature inputs for each node in graph convolutional networks (GCN). The specific network structure is expressed as follows:
@@ 4.3 Transformer-Based Multi-key Gated Communication Mechanism
After constructing $S(o^{i}_{t},m^{i}_{t},A^{i}_{t})$, we designed a multi-key gated mechanism to learn $A^{i}_{t}$ in the Q-network. Combined with the transformer decoder, it primarily addresses the first four of the five communication problems discussed in Section 1. The multi-key gated communication network is the encoder, and the transformer is the decoder. The schematics of the communication mechanism are shown in Figure 2(c).
@@
where gumbel-softmax (Jang, Gu, and Poole 2016) is an zactivation function for two output dimensions (communicate or not). After passing through communication network, the adjacency trajectory matrix is expressed as follows:
@@ 5 Experiments
In this section, we evaluate our approach against five representative MARL baselines on various benchmarks with differing communication requirements. QMIX (Rashid et al. 2020) is a strong non-communication baseline, showing outstanding performance across multiple multi-agent benchmarks(Papoudakis et al. 2020). IC3Net (Singh, Jain, and Sukhbaatar 2018) includes a gating mechanism for learning when to broadcast messages, using CommNet (Sukhbaatar, Fergus et al. 2016). DGN (Jiang et al. 2018) uses graph convolutional networks with relational kernels to capture dynamic agent interactions. G2ANet (Liu et al. 2020) builds a sparse communication interaction graph with a two-stage attention mechanism. GraphMix (Naderializadeh et al. 2020) also constructs a communication model using graph structures and is a notably effective baseline. All algorithms use the EPyMARL (Papoudakis et al. 2020) implementation. We evaluate TGCNet with multiple state-of-the-art baselines on tasks including Hallway (Wang et al. 2019), level-based foraging (LBF) (Papoudakis et al. 2020), and the SMAC (Samvelyan et al. 2019). Figure 3(a) shows the Hallway task, where four agents start at different states ($a_{1}$ to $a_{x}$, $b_{1}$ to $b_{y}$, and $c_{1}$ to $c_{z}$, with $x,y,z=4,6,10$) and must reach the goal $g$. Figure 3(b) depicts the LBF task, where six agents cooperate to collect four portions of food in an $11\times 1$ 11 grid world. In SMAC, the 1o2r_vs_4r and 1o10b_vs_1r maps (Wang et al. 2019) require collaboration and communication among agents to identify enemy positions, as shown in Figure 3(c). For fair evaluation, all experiments are conducted with five random seeds, and results are presented as means with a 95% confidence interval. Detailed information on benchmarks and hyper-parameter selections is provided in the Appendix.
@@ 5.1 Overall Performance Comparison
We started by evaluating the overall performance of TGCNet against multiple baselines in communication-intensive benchmarks. As shown in Figure 4, TGCNet consistently outperforms the other methods with low variance across all benchmarks, demonstrating its robustness in scenarios of varying difficulty. In the Hallway task (Figure 4(a)), which requires frequent communication for success, methods lacking communication capabilities, such as QMIX, fail. Other communication-based methods also perform poorly or fail entirely in this environment. This finding indicates that inappropriate message generation or selection would injure the learning process. The performance of TGCNet exceeds GraphMix by approximately 20% given its effective communication modeling and powerful feature extraction capabilities. In the LBF task (Figure 4(b)), existing communication-based MARL methods, such as IC3Net, G2ANet, and DGN encounter challenges due to sparse rewards, especially when the food items are widely distributed. Different from its performance in Hallway, QMIX performs effectively in LBF because agents can observe nearby grids. Our method matches the performance of GraphMix, demonstrating strong coordination abilities even in scenarios with sparse rewards. For scenarios requiring communication in SMAC, QMIX performs the worst without a communication mechanism, whereas TGCNet can maintain high efficiency of learning and exhibit consistently competitive performance when converged, superior to other baselines.
To validate our algorithm’s information extraction and scalability, we tested it in SMAC scenarios without communication, including hard and super hard levels. As shown in Figure 5, TGCNet achieves the highest average test win rate across all scenarios upon convergence. GraphMix performs slightly worse than TGCNet but has a faster convergence speed and higher win rate than DGN and G2ANet due to its graph-based communication information aggregation. IC3Net, DGN, and G2ANet perform worse than QMIX in scenarios without communication because their inaccurate communication modeling fails to manage redundant information effectively, impacting the training of reinforcement learning policy networks. In hard scenarios, TGCNet’s average test win rate is nearly 20% higher than those of algorithms. This finding is attributed to the transformer-based decoder’s superior representation of policy networks. In super hard scenarios with more agents, especially on the 27m_vs_30m map, TGCNet maintains a higher convergence speed due to its multi-key gated communication network. Overall, TGCNet demonstrates excellent performance in all scenarios, learning effective communication strategies through its transformer-based multi-key gated communication mechanism and accurately adjusting the global state with the graph coarsening network.
@@ 5.2 Ablation Studies
To understand the superior performance of TGCNet, we conducted ablation studies to evaluate the contributions of its two main components, addressing the following questions: (1) Is the graph coarsening network effective for fitting the global state? How does it compare to using real global states? (2) How does the algorithm’s effectiveness differ between full and partial communication? To address the first question, we created the TGCNet_QMix algorithm, replacing TGCNet’s graph coarsening network with a hybrid network using QMIX, while retaining other network structures. For the second question, we created TGCNet_FC, which removes TGCNet’s multi-key communication network and uses full communication (FC). We conducted ablation experiments on one hard scenario and two super hard scenarios. The results shown in Figure 6 illustrate that TGCNet performs similarly to TGCNet_QMIX in these scenarios, indicating that TGCNet’s graph coarsening network effectively approximates the global state. TGCNet_FC converges more gradually than TGCNet in the corridor scenario but ultimately achieves the same performance level. This finding suggests that the multi-key communication network optimizes the communication structure without sacrificing performance, thereby avoiding communication waste.
@@ 5.3 Communication Performance
To analyze the communication strategies learned through the graph structure, we conducted a visualization analysis of the multi-agent communication network after training, focusing on a test scenario in SMAC. We illustrate the evolution of the graph structure over the trajectory, as shown in the Figure 7. Initially, at $t=0$, no enemies were found within the allies’ field of view, and the agents were unaware of each other’s information. This condition resulted in a fully connected graph, indicating a state of complete communication. In the next timestep, if the agents’ observations and states remained unchanged after their initial communication, the graph structure connections and communication volume decreased. At $t=3$, communication increased again when enemies entered the agents’ fields of view. Similarly, in subsequent timesteps, minimal information changes led to a synchronous reduction in communication. In summary, the visualization results indicate that the multi-key communication network learned by TGCNet possesses a certain level of interpretability.

### v2

@@ 4.2 Graph Coarsening Network
We construct the graph coarsening network based on Equation (2) to address the last of the five communication problems outlined in Section 1. The schematics of the graph coarsening network are shown in Figure 2(a). The dynamic directed graph, composed of multiple subgraphs, can be considered a coarsened graph, where each subgraph is a super node to fit the global state. This approach enables hierarchical learning of global information. According to the theoretical analysis in Section 4.1, the input of the graph coarsening network consists of each agent’s local observations $o^{i}_{t}$, the adjacency trajectory matrix $A^{i}_{t}$ and the received message $\tilde{m}^{i}_{t}$. We normalize the adjacency trajectory matrix to prevent degree bias and gradient vanishing during training. We define $[\tilde{m}^{i}_{t}\parallel o^{i}_{t}]$ as $X^{i}_{t}$ as feature inputs for each node in graph convolutional networks (GCN). The specific network structure is expressed as follows:
@@ 4.3 Transformer-Based Multi-Key Gated Communication Mechanism
After constructing $S(o^{i}_{t},m^{i}_{t},A^{i}_{t})$, we design a multi-key gated mechanism to learn $A^{i}_{t}$ in the Q-network. Combined with the Transformer decoder, it addresses the first four of the five communication problems discussed in Section 1. The multi-key gated communication network is the encoder, and the Transformer is the decoder. The schematics of the communication mechanism are shown in Figure 2(c).
@@
where gumbel-softmax (Jang, Gu, and Poole 2016) is an activation function for two output dimensions (communicate or not). After passing through communication network, the adjacency trajectory matrix is expressed as follows:
@@ 5 Experiments
In this section, we evaluate our approach against five representative MARL baselines on various benchmarks with differing communication requirements. QMIX (Rashid et al. 2020) is a strong non-communication baseline, showing outstanding performance across multiple multi-agent benchmarks(Papoudakis et al. 2020). IC3Net (Singh, Jain, and Sukhbaatar 2018) includes a gating mechanism for learning when to broadcast messages, using CommNet (Sukhbaatar, Fergus et al. 2016). DGN (Jiang et al. 2018) uses graph convolutional networks with relational kernels to capture dynamic agent interactions. G2ANet (Liu et al. 2020) builds a sparse communication interaction graph with a two-stage attention mechanism. GraphMix (Naderializadeh et al. 2020) also constructs a communication model using graph structures and is a notably effective baseline. All algorithms use the EPyMARL (Papoudakis et al. 2020) implementation. We evaluate TGCNet with multiple state-of-the-art baselines on tasks including Hallway (Wang et al. 2019), LBF (Papoudakis et al. 2020), and the SMAC (Samvelyan et al. 2019). Figure 3(a) shows the Hallway task, where four agents start at different states ($a_{1}$ to $a_{x}$, $b_{1}$ to $b_{y}$, and $c_{1}$ to $c_{z}$, with $x,y,z=4,6,10$) and must reach the goal $g$. Figure 3(b) depicts the LBF task, where six agents cooperate to collect four portions of food in an $11\times 1$ 11 grid world. In SMAC, the 1o2r_vs_4r and 1o10b_vs_1r maps (Wang et al. 2019) require collaboration and communication among agents to identify enemy positions, as shown in Figure 3(c). For fair evaluation, all experiments are conducted with five random seeds, and results are presented as means with a 95% confidence interval.
@@ 5.1 Overall Performance Comparison
We start by evaluating the overall performance of TGCNet against multiple baselines in communication-intensive benchmarks. As shown in Figure 4, TGCNet consistently outperforms the other methods with low variance across all benchmarks, demonstrating its robustness in scenarios of varying difficulty. In the Hallway task (Figure 4(a)), which requires frequent communication for success, methods lacking communication capabilities, such as QMIX, fail. Other communication-based methods also perform poorly or fail entirely in this environment. This finding indicates that inappropriate message generation or selection would injure the learning process. The performance of TGCNet exceeds GraphMix by approximately 20% given its effective communication modeling and powerful feature extraction capabilities. In the LBF task (Figure 4(b)), existing communication-based MARL methods, such as IC3Net, G2ANet, and DGN encounter challenges due to sparse rewards, especially when the food items are widely distributed. Different from its performance in Hallway, QMIX performs effectively in LBF because agents can observe nearby grids. Our method matches the performance of GraphMix, demonstrating strong coordination abilities even in scenarios with sparse rewards. For scenarios requiring communication in SMAC, QMIX performs the worst without a communication mechanism, whereas TGCNet can maintain high efficiency of learning and exhibit consistently competitive performance when converged, superior to other baselines.
To validate our algorithm’s information extraction and scalability, we test it in SMAC scenarios without communication, including hard and super hard levels. As shown in Figure 5, TGCNet achieves the highest average test win rate across all scenarios upon convergence. GraphMix performs slightly worse than TGCNet but has a faster convergence speed and higher win rate than DGN and G2ANet due to its graph-based communication information aggregation. IC3Net, DGN, and G2ANet perform worse than QMIX in scenarios without communication because their inaccurate communication modeling fails to manage redundant information effectively, impacting the training of reinforcement learning policy networks. In hard scenarios, TGCNet’s average test win rate is nearly 20% higher than those of algorithms. This finding is attributed to the Transformer-based decoder’s superior representation of policy networks. In super hard scenarios with more agents, especially on the 27m_vs_30m map, TGCNet maintains a higher convergence speed due to its multi-key gated communication network. Overall, TGCNet demonstrates excellent performance in all scenarios, learning effective communication strategies through its Transformer-based multi-key gated communication mechanism and accurately adjusting the global state with the graph coarsening network.
@@ 5.2 Ablation Studies
To understand the superior performance of TGCNet, we conduct ablation studies to evaluate the contributions of its two main components, addressing the following questions: (1) Is the graph coarsening network effective for fitting the global state? How does it compare to using real global states? (2) How does the algorithm’s effectiveness differ between full and partial communication? To address the first question, we create the TGCNet_QMix algorithm, replacing TGCNet’s graph coarsening network with a hybrid network using QMIX, while retaining other network structures. For the second question, we create TGCNet_FC, which removes TGCNet’s multi-key communication network and uses full communication (FC). We conduct ablation experiments on one hard scenario and two super hard scenarios. The results shown in Figure 6 illustrate that TGCNet performs similarly to TGCNet_QMIX in these scenarios, indicating that TGCNet’s graph coarsening network effectively approximates the global state. TGCNet_FC converges more gradually than TGCNet in the corridor scenario but ultimately achieves the same performance level. This finding suggests that the multi-key communication network optimizes the communication structure without sacrificing performance, thereby avoiding communication waste.
@@ 5.3 Communication Performance
To analyze the communication strategies learned through the graph structure, we conduct a visualization analysis of the multi-agent communication network after training, focusing on a test scenario in SMAC. We illustrate the evolution of the graph structure over the trajectory, as shown in the Figure 7. Initially, at $t=0$, no enemies are found within the allies’ field of view, and the agents are unaware of each other’s information. This condition result in a fully connected graph, indicating a state of complete communication. In the next timestep, if the agents’ observations and states remain unchanged after their initial communication, the graph structure connections and communication volume decreased. At $t=3$, communication increas again when enemies entered the agents’ fields of view. Similarly, in subsequent timesteps, minimal information changes lead to a synchronous reduction in communication. In summary, the visualization results indicate that the multi-key communication network learned by TGCNet possesses a certain level of interpretability.
@@ 6 Conclusions and Future Work
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under grants 62103302 and U1713215, by Shanghai Rising-Star Program under grant 24QA2709400, by the Shanghai Chenguang Program under grant 22CGA19, by the Shanghai Municipal Science and Technology Major Project under grant 2021SHZDZX0100, and by Fundamental Research Funds for the Central Universities under grant 22120240276.